A small factory producing a single product has weekly fixed costs of production of $2,112 and weekly variable costs of $52x + 3/4 x2, where x is the quantity produced. the capacity of the factory is about 600 units.


Past experience suggests that the product’s price and quantity are linked by the following demand equation: p = 200 - 1/4 x (p, x > 0) where p = $ price/unit and x = quantity sold. You are required to:


(a) Find the level of production at which revenue is maximized


(b) Find any break-even points

1
Expert's answer
2022-01-18T18:14:33-0500

(a) The level of production at which revenue is maximized is:

"MR = TR'(x) = (p\u00d7x)' = 200 - 0.5x = 0,"

0.5x = 200,

x = 400 units.

(b) The break-even points are:

"x = FC\/(p - AVC) = 2,112\/(200 - 1\/4x - 52 - 3\/4x) =2,112\/(148 - x),"

"x^2 - 148x + 2,112 = 0,"

"x = (148 \pm 116)/2,"

"x_1 = 132" units, "x_2 = 16" units.